# Lyapunov exponents and geodesic stability of Schwarzschild black hole in the non-commutative gauge theory of gravity

### v3

@@ III.2 Circular motion of massive particles and Lyapunov exponents
Fig. 4 illustrates the influence of the NC parameter $\Theta$ on the effective potential of a massive test particle. Notably, the introduction of non-commutativity generates a new minimum in the effective potential near the event horizon, regardless of the angular momentum values ($l\geq 0$). This new minimum, located outside the event horizon, corresponds to a stable circular orbit abdellah1 , leading to the existence of multiple stable circular orbits in NC Schwarzschild geometry. In contrast, prior studies on non-commutativity in geodesic motion—whether based on NC matter distributions nozari1 ; nozari2 ; kuniyal ; iftikhar ; larranaga ; bhar ; rahaman ; NCGD1 ; NCGD3 ; NCLD2 ; NCLD3 and on NC gauge theory of gravity via deformed geometry NCGTG2 ; NCGTG3 or via deformed mass NCGTGmassdeformed1 ; NCGTGmassdeformed2 ; NCGTGmassdeformed3 have not predicted these new stable circular orbits near the event horizon.
@@
$E^{2}_{c}$ $\simeq\frac{(-2m+r_{c})^{2}}{r_{c}(r_{c}-3m)}-\frac{\left(B(r_{c})+F(r_{c})\sqrt{1-\frac{2m}{r_{c}}}\right)\Theta^{2}}{32r_{c}^{4}(-3m+r_{c})^{2}\sqrt{1-\frac{2m}{r_{c}}}}+\mathcal{O}(\Theta^{4}),$ (34)
$L^{2}_{c}$ $\simeq\frac{mr_{c}^{2}}{r_{c}-3m}-\frac{\left(S(r_{c})+Q(r_{c})\sqrt{1-\frac{2m}{r_{c}}}\right)\Theta^{2}}{32r_{c}(-3m+r_{c})^{2}(r_{c}-2m)\sqrt{1-\frac{2m}{r_{c}}}}+\mathcal{O}(\Theta^{4}),$ (35)
@@ III.2.1 Orbital velocity and time period
$\hat{\Omega}_{c}=\sqrt{\frac{m}{r_{c}^{3}}+\frac{X(r_{c})+W(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{32r_{c}^{6}(r_{c}-2m)^{3}}\Theta^{2}},$ (38)
@@
Fig. 6 illustrates the orbital velocity (left panel) and the orbital time scale (right panel) of a massive test particle in NC Schwarzschild spacetime as functions of $r_{c}$. The effects of non-commutativity become significant near the NC event horizon $r_{h}^{NC}$. Specifically, non-commutativity increases the orbital velocity in the vicinity of $r_{h}^{NC}$, while its impact become negligible at larger orbital radius. Similarly, the orbital period decreases due to non-commutativity, approaching zero near the NC event horizon. This reduction is most pronounced close to $r_{h}^{NC}$ and becomes negligible for $r_{c}\gg 3$, where NC effects vanish.
@@ III.2.2 Lyapunov exponents
To analyze the stability and instability of circular geodesics for a massive test particle in the NC Schwarzschild spacetime, we employ the Lyapunov exponents $\hat{\lambda}$. The stability of an orbit is determined by the nature of $\hat{\lambda}$: an imaginary $\hat{\lambda}$ indicates a stable orbit, whereas a real $\hat{\lambda}$ signifies instability geostab2 ; geostab10 . The proper time ($\hat{\lambda}_{p}$) and coordinate time ($\hat{\lambda}_{c}$) Lyapunov exponents in NC spacetime are given by geostab3 ; geostab2 ; geostab10 ; geostab1 :
@@
$\hat{\lambda}_{p}$ $=\sqrt{-\frac{m(r_{c}-6m)}{(r_{c}-3m)r_{c}^{3}}+\frac{Z(r_{c})+P(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{32(r_{c}-2m)^{3}r_{c}^{6}(r_{c}-3m)^{2}}\Theta^{2}},$ (42)
$\hat{\lambda}_{c}$ $=\sqrt{-\frac{m(r_{c}-6m)}{r_{c}^{4}}+\frac{Y(r_{c})+N(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{32(r_{c}-2m)^{3}r_{c}^{7}}\Theta^{2}}.$ (43)
@@
Fig. 7 illustrates the proper time (left panel) and coordinate time (right panel) Lyapunov exponents for a massive test particle in NC spacetime. The coordinate time Lyapunov exponent $\hat{\lambda}_{c}$ is real in the region $r_{h}^{NC}<r_{c}\leq r_{outer}^{ISCO}$, indicating unstable orbits near the event horizon. As the orbit radius $r_{c}$ increases, the instability decreases, and reaching zero at the outer ISCO, $r_{outer}^{ISCO}$. The proper time Lyapunov exponent $\hat{\lambda}_{p}$ exhibits a distinct behavior: instability (real values) is observed in two regions, $r_{h}^{NC}<r_{c}\leq r_{inner}^{ISCO}$ and $r_{c}^{uns}<r_{c}\leq r_{outer}^{ISCO}$. Between these intervals, in the range $r_{inner}^{ISCO}\leq r_{c}\leq r_{c}^{uns}$, $\hat{\lambda}_{p}$ becomes imaginary, indicating stability of the circular orbit. These results confirm the existence of new stable circular orbits near the event horizon, consistent with previous findings abdellah1 .
The ratio of the proper time to coordinate time Lyapunov exponents is given by:
$\frac{\hat{\lambda}_{p}}{\hat{\lambda}_{c}}=\sqrt{\frac{r_{c}}{r_{c}-3m}+\frac{Y(r_{c})(3m-r_{c})-Z(r_{c})+(N(r_{c})(3m-r_{c})-P(r_{c}))\sqrt{1-\frac{2m}{r_{c}}}}{32m(r_{c}-2m)^{3}(r_{c}-3m)^{2}(r_{c}-6m)r_{c}^{2}}\Theta^{2}}.$ (45)
@@
Fig. 8 illustrates the ratio $\hat{\lambda}_{p}/\hat{\lambda}_{c}$ as a function of $r_{c}$ in NC spacetime. The ratio remains real in two distinct regions, separated by an interval where it becomes imaginary, reflecting the behavior observed in the proper time Lyapunov exponent. The analysis of proper time and coordinate time Lyapunov exponents, along with their ratio, reveals a novel feature in the dynamics of circular motion around an NC Schwarzschild BH. Specifically, it predicted a new range of stable circular orbits located between unstable ones.
@@ III.3 Radial motion of massless particles
The radial motion of massless particles in the NC Schwarzschild spacetime can be derived using the effective potential (23) by setting $h=0$, $c=0$, and $l=0$. This leads to the equation governing radial motion:
$\left(\frac{dr}{d\tau}\right)^{2}$ $=E^{2}-\Theta^{2}\left\{E^{2}\frac{m\big{(}64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r}})r+2(13-3\sqrt{1-\frac{2m}{r}})r^{2}\big{)}}{16r^{5}\left(1-\frac{2m}{r}\right)^{2}}\right\}+\mathcal{O}(\Theta^{4}),$ (46)
where the commutative case is recovered when $\Theta=0$.
For a photon freely falling toward the NC Schwarzschild BH and emitted from a radial position $r=r_{0}$ with energy $E=1$, the affine parameter $\tau$ and coordinate time $\hat{t}$ are given by:
@@
$\hat{\tau}$ $=r_{0}-r+\Theta^{2}\bigg{\{}25\ln\left(\frac{r_{0}(r-2m)}{r(r_{0}-2m)}\right)-\frac{1}{128m}\left[\frac{32m^{2}+15mr_{0}+8r_{0}^{2}}{r_{0}^{2}}-\frac{32m^{2}+15mr+8r^{2}}{r^{2}}\right]$
$\quad+\frac{1}{128m}\bigg{[}\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}-\left(1-\frac{2m}{r_{0}}\right)^{-\frac{1}{2}}-14\big{(}\sqrt{1-\frac{2m}{r}}-\sqrt{1-\frac{2m}{r_{0}}}\big{)}$
$\quad-\frac{13}{3}\big{(}(1-\frac{2m}{r})^{\frac{3}{2}}-(1-\frac{2m}{r_{0}})^{\frac{3}{2}}\big{)}-\frac{35}{2}\frac{r_{0}(r-2m)-r(r_{0}-2m)}{(r_{0}-2m)(r-2m)}\bigg{]}\bigg{\}},$ (49)
$\hat{t}$ $=r_{0}-r+2m\ln\left(\frac{r_{0}-2m}{r-2m}\right)+\Theta^{2}\bigg{\{}-\frac{224m^{2}-201mr_{0}+63r_{0}^{2}}{128r_{0}(r_{0}-2m)^{2}}+\frac{224m^{2}-201mr+63r^{2}}{128r(r-2m)^{2}}$
$\quad-\frac{(17m^{2}-24mr_{0}+8r_{0}^{2})\sqrt{1-\frac{2m}{r_{0}}}}{m(r_{0}-2m)^{2}}+\frac{(17m^{2}-24mr+8r^{2})\sqrt{1-\frac{2m}{r}}}{m(r-2m)^{2}}-\frac{63}{m}\ln\left(\frac{r_{0}(r-2m)}{r(r_{0}-2m)}\right)\bigg{\}}.$ (50)
@@
Fig. 9 illustrates the behavior of $\hat{\tau}$ and $\hat{t}$ for a photon initially at $r=r_{0}$ falling toward the NC Schwarzschild BH. Notably, in the affine parameter framework of NC spacetime, photons require an infinite affine parameter to reach the NC singularity, in contrast to the commutative case Chandr1 ; radial1 . The same observation applies to coordinate time, where the photon cannot reach the singularity within a finite time. This behavior is consistent with the motion of massive particles and highlights the role of non-commutativity in preventing singularity encounters within a finite temporal or affine framework. Additionally, for massless particles, our analysis predicts a novel behavior in free-fall motion within NC spacetime using the NC gauge theory of gravity. This differs from the approach in Refs. rahaman ; bhar , which alters the commutative behavior and modifies the underlying physics of the spacetime geometry. In contrast, non-commutativity should act as a quantum correction to the metric, reducing to the commutative case when this correction vanishes. While using an NC matter distribution can recover the commutative free-fall motion, as demonstrated in Ref. kuniyal , it does not introduce any new predictions for free-fall dynamics. However, our approach based on the NC gauge theory of gravity preserves the commutative solution while predicting new macroscopic effects of non-commutativity.
III.4 Circular motion of massless particles and Lyapunov exponents
This section examines the null geodesic motion of photons around an NC Schwarzschild black hole (BH). The effective potential for a photon in this spacetime is obtained from (23) by setting $h=0$, $c=1$ and $l\neq 0$.
@@ III.4 Circular motion of massless particles and Lyapunov exponents
Fig. 10 illustrates the influence of the NC parameter $\Theta$ (left panel) and the angular momentum $l$ (right panel) on the effective potential of a massless test particle (photon). As shown in the left panel, introducing non-commutativity leads to the emergence of a new minimum in the effective potential close to the event horizon. This phenomenon is analogous to what is observed for massive test particles (see SubSec. III.2). However, for massless particles, only two extrema of the effective potential exist outside the event horizon, allowing the interpretation of the new minimum as a stable circular orbit. This results in the formation of a stable photon sphere in this spacetime. Additionally, the position of this new minimum shifts farther from the event horizon as $\Theta$ increases. Moreover, the influence of non-commutativity becomes negligible in regions significantly distant from the event horizon ($r\gg r_{h}^{NC}$). The right panel of Fig. 10 reveals that in NC spacetime, the effective potential exhibits extremal points only for angular momentum values that satisfy a specific condition. This contrasts with the commutative case, where the effective potential always possesses a maximum, irrespective of the angular momentum. For $\Theta=0.2$, the effective potential exhibits two extrema when the angular momentum exceeds a critical value, $l^{\text{crit}}\geq 2.55$. These critical values depend on the NC parameter $\Theta$. In conclusion, NC spacetime consistently supports the existence of two photon spheres: an inner stable photon sphere and an outer unstable photon sphere. This behavior differs significantly from the commutative case and highlights the unique role of NC geometry in modifying the geodesic structure of spacetime. Furthermore, these results are novel and distinct from previous works kuniyal ; bhar ; rahaman ; NCLD2 ; NCGTG2 ; NCGTG3 ; NCGTGmassdeformed1 ; NCGTGmassdeformed2 ; NCGTGmassdeformed3 , which could not predict the existence of a new stable photon sphere near the event horizon.

### v4

@@ III.2 Circular motion of massive particles
Fig. 5 illustrates the influence of the NC parameter $\Theta$ on the effective potential of a massive test particle. Notably, the introduction of non-commutativity generates a new minimum in the effective potential near the event horizon, regardless of the angular momentum values ($l\geq 0$). This new minimum, located outside the event horizon, corresponds to a stable circular orbit abdellah1 , leading to the existence of multiple stable circular orbits in NC Schwarzschild geometry.
In contrast, prior studies on non-commutativity in geodesic motion whether based on NC matter distributions nozari1 ; nozari2 ; kuniyal ; iftikhar ; larranaga ; bhar ; rahaman ; NCGD1 ; NCGD3 ; NCLD2 ; NCLD3 , NC gauge theory of gravity via deformed geometry NCGTG2 ; NCGTG3 , or deformed mass NCGTGmassdeformed1 ; NCGTGmassdeformed2 ; NCGTGmassdeformed3 have not predicted these new stable circular orbits near the event horizon.
@@
$E^{2}_{c}$ $\simeq\frac{(r_{c}-2m)^{2}}{r_{c}(r_{c}-3m)}-\frac{\left(B(r_{c})+F(r_{c})\sqrt{1-\frac{2m}{r_{c}}}\right)\Theta^{2}}{8r_{c}^{4}(-3m+r_{c})^{2}\sqrt{1-\frac{2m}{r_{c}}}}+\mathcal{O}(\Theta^{4}),$ (33)
$L^{2}_{c}$ $\simeq\frac{mr_{c}^{2}}{r_{c}-3m}-\frac{\left(S(r_{c})+Q(r_{c})\sqrt{1-\frac{2m}{r_{c}}}\right)\Theta^{2}}{8r_{c}(r_{c}-3m)^{2}(r_{c}-2m)\sqrt{1-\frac{2m}{r_{c}}}}+\mathcal{O}(\Theta^{4}),$ (34)
@@ III.2.1 Orbital velocity and time period
$\hat{\Omega}_{c}=\sqrt{\frac{m}{r_{c}^{3}}+\frac{X(r_{c})+W(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{8r_{c}^{6}(r_{c}-2m)^{3}}\Theta^{2}},$ (37)
@@
Fig. 7 illustrates the orbital velocity (left panel) and the orbital time scale (right panel) of a massive test particle in NC Schwarzschild spacetime as functions of $r_{c}$. The effects of non-commutativity become significant near the NC event horizon $r_{h}^{NC}$. Specifically, non-commutativity increases the orbital velocity in the vicinity of $r_{h}^{NC}$, while its impact becomes negligible at larger orbital radii. Similarly, the orbital period decreases due to non-commutativity, approaching zero near the NC event horizon. This reduction is most pronounced close to $r_{h}^{NC}$ and becomes negligible for $r_{c}\gg 3$, where NC effects vanish.
@@ III.2.2 Lyapunov exponents
To analyze the stability and instability of circular geodesics for a massive test particle in the NC Schwarzschild spacetime, we employ the Lyapunov exponents $\hat{\lambda}$. The stability of an orbit is determined by the nature of $\hat{\lambda}$: an imaginary $\hat{\lambda}$ indicates a stable orbit, whereas a real $\hat{\lambda}$ signifies instability geostab2 ; geostab10 . The proper-time ($\hat{\lambda}_{p}$) and coordinate-time ($\hat{\lambda}_{c}$) Lyapunov exponents in NC spacetime are given by geostab3 ; geostab2 ; geostab10 ; geostab1 :
@@
$\hat{\lambda}_{p}$ $=\sqrt{-\frac{m(r_{c}-6m)}{(r_{c}-3m)r_{c}^{3}}+\frac{Z(r_{c})+P(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{16(r_{c}-2m)^{3}r_{c}^{6}(r_{c}-3m)^{2}}\Theta^{2}},$ (41)
$\hat{\lambda}_{c}$ $=\sqrt{-\frac{m(r_{c}-6m)}{r_{c}^{4}}+\frac{Y(r_{c})+N(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{16(r_{c}-2m)^{3}r_{c}^{7}}\Theta^{2}}.$ (42)
@@
Fig. 8 illustrates the proper-time (left panel) and coordinate-time (right panel) Lyapunov exponents for a massive test particle in NC spacetime. The coordinate-time Lyapunov exponent $\hat{\lambda}_{c}$ is real in the region $r_{h}^{NC}<r_{c}\leq r_{outer}^{ISCO}$, indicating unstable orbits near the event horizon. As the orbital radius $r_{c}$ increases, the instability decreases, reaching zero at the outer ISCO, $r_{outer}^{ISCO}$. The proper-time Lyapunov exponent $\hat{\lambda}_{p}$ exhibits distinct behavior: instability (real values) appears in two regions, $r_{h}^{NC}<r_{c}\leq r_{inner}^{ISCO}$ and $r_{c}^{uns}<r_{c}\leq r_{outer}^{ISCO}$. However, in the intermediate range $r_{inner}^{ISCO}\leq r_{c}\leq r_{c}^{uns}$, $\hat{\lambda}_{p}$ becomes imaginary, indicating stable circular orbits. These findings confirm the existence of new stable circular orbits near the event horizon, consistent with previous results abdellah1 .
The ratio of the proper-time to coordinate-time Lyapunov exponents is given by:
$\frac{\hat{\lambda}_{p}}{\hat{\lambda}_{c}}=\sqrt{\frac{r_{c}}{r_{c}-3m}+\frac{Y(r_{c})(3m-r_{c})-Z(r_{c})+(N(r_{c})(3m-r_{c})-P(r_{c}))\sqrt{1-\frac{2m}{r_{c}}}}{16m(r_{c}-2m)^{3}(r_{c}-3m)^{2}(r_{c}-6m)r_{c}^{2}}\Theta^{2}}.$ (44)
@@
Fig. 9 illustrates the ratio $\hat{\lambda}_{p}/\hat{\lambda}_{c}$ as a function of $r_{c}$ in NC spacetime. This ratio remains real in two distinct regions, separated by an interval where it becomes imaginary, mirroring the behavior observed in the proper-time Lyapunov exponent. The analysis of proper-time and coordinate-time Lyapunov exponents, along with their ratio, reveals a novel feature in the dynamics of circular motion around an NC Schwarzschild BH: the existence of a new range of stable circular orbits interspersed between unstable ones.
@@ III.3 Radial motion of massless particles
In this subsection, we study the free-fall motion of a massless test particle in this geometry. The motion is described by the radial trajectory of a massless particle in the NC Schwarzschild spacetime and is obtained using the effective potential (22), setting $h=0$, $c=0$, and $l=0$. This leads to the equation governing radial motion:
$\left(\frac{dr}{d\tau}\right)^{2}$ $=E^{2}-\Theta^{2}\left\{E^{2}\left(\frac{m\left(128m^{2}-12r^{2}\left(\sqrt{1-\frac{2m}{r}}-3\right)+mr\left(26\sqrt{1-\frac{2m}{r}}-99\right)\right)}{16r^{3}(r-2m)^{2}}\right)\right\}+\mathcal{O}(\Theta^{4}),$ (45)
where the commutative case is recovered for $\Theta=0$. For a photon freely falling toward the NC Schwarzschild BH, emitted from a radial position $r=r_{0}$ with energy $E=1$, the affine parameter $\tau$ and coordinate time $\hat{t}$ are given by:
@@
$\hat{\tau}$ $=r_{0}-r+\frac{\Theta^{2}}{128m}\bigg{\{}33\ln\left(\frac{r_{0}(r-2m)}{r(r_{0}-2m)}\right)-\left[\frac{m(64m+29r_{0})+16r_{0}^{2}}{r_{0}^{2}}-\frac{m(64m+29r)+16r^{2}}{r^{2}}\right]$
$\quad+\bigg{[}2\Bigg{(}\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}-\left(1-\frac{2m}{r_{0}}\right)^{-\frac{1}{2}}\Bigg{)}+28\bigg{(}\sqrt{1-\frac{2m}{r}}-\sqrt{1-\frac{2m}{r_{0}}}\bigg{)}$
$\quad+\frac{26}{3}\bigg{(}\bigg{(}1-\frac{2m}{r_{0}}\bigg{)}^{\frac{3}{2}}-\bigg{(}1-\frac{2m}{r}\bigg{)}^{\frac{3}{2}}\bigg{)}-\frac{37}{2}\frac{r_{0}(r-2m)-r(r_{0}-2m)}{(r_{0}-2m)(r-2m)}\bigg{]}\bigg{\}},$ (48)
$\hat{t}$ $=r_{0}-r+2m\ln\left(\frac{r_{0}-2m}{r-2m}\right)+\frac{\Theta^{2}}{128m}\bigg{\{}\frac{r^{2}(104m-11r)+448m(r-2m)^{2}}{4r(r-2m)^{2}}-\frac{r_{0}^{2}(104m-11r_{0})+448m(r_{0}-2m)^{2}}{4r_{0}(r_{0}-2m)^{2}}$
$\quad+8\Bigg{(}\frac{(17m^{2}-24mr+8r^{2})\sqrt{1-\frac{2m}{r}}}{m(r-2m)^{2}}-\frac{(17m^{2}-24mr_{0}+8r_{0}^{2})\sqrt{1-\frac{2m}{r_{0}}}}{m(r_{0}-2m)^{2}}\Bigg{)}-\frac{127}{2}\ln\left(\frac{r_{0}(r-2m)}{r(r_{0}-2m)}\right)\bigg{\}}.$ (49)
@@
Fig. 10 illustrates the behavior of $\hat{\tau}$ and $\hat{t}$ for a photon initially at $r=r_{0}$ falling toward the NC Schwarzschild BH. Notably, in the affine parameter framework of NC spacetime, photons require an infinite affine parameter to reach the NC singularity, in contrast to the commutative case Chandr1 . The same observation applies to coordinate time, where the photon cannot reach the singularity within a finite time. This behavior aligns with the motion of massive particles and highlights the role of non-commutativity in preventing singularity encounters within a finite temporal or affine framework. Additionally, for massless particles, our analysis predicts a novel free-fall behavior in NC spacetime using the NC gauge theory of gravity. This differs from the NC matter distribution approach in Refs. rahaman ; bhar , which modifies the commutative behavior and alters the underlying physics of the spacetime geometry. In contrast, non-commutativity should act as a quantum correction to the metric, reducing to the commutative case when this correction vanishes. While an NC matter distribution can recover the commutative free-fall motion, as demonstrated in Ref. kuniyal , it does not introduce new predictions for free-fall dynamics. However, our approach, based on the NC gauge theory of gravity, preserves the commutative solution while predicting novel macroscopic effects of non-commutativity.
III.4 Circular motion of massless particles
This section explores the null geodesic motion of photons in the NC Schwarzschild black hole (BH) spacetime. The effective potential for a photon is derived from (22) by setting $h=0$, $c=1$, and $l\neq 0$.
@@ III.4 Circular motion of massless particles
Fig. 11 illustrates the influence of the NC parameter $\Theta$ (left panel) and the angular momentum $l$ (right panel) on the effective potential of a massless test particle (photon). As shown in the left panel, introducing non-commutativity leads to the emergence of a new minimum in the effective potential near the event horizon. This phenomenon is analogous to that observed for massive test particles (see SubSec. III.2). However, for massless particles, only two extrema exist outside the event horizon, allowing the interpretation of the new minimum as a stable circular orbit. Consequently, this results in the formation of a stable photon sphere in NC spacetime. Additionally, the position of this new minimum shifts farther from the event horizon as $\Theta$ increases, while NC effects become negligible at large distances ($r\gg r_{h}^{NC}$). The right panel of Fig. 11 reveals that, in NC spacetime, the effective potential exhibits extremal points only when the angular momentum satisfies a specific condition. This contrasts with the commutative case, where the effective potential always possesses a maximum, regardless of the angular momentum. For $\Theta=0.2$, the effective potential exhibits two extrema when the angular momentum exceeds a critical value, $l^{\text{crit}}\geq 2.55$. These critical values depend on the NC parameter $\Theta$. In conclusion, NC spacetime consistently supports the existence of two photon spheres: an inner stable photon sphere and an outer unstable photon sphere. This behavior significantly differs from the commutative case and highlights the unique role of NC geometry in modifying the geodesic structure of spacetime. Furthermore, these results are novel and distinct from previous works using different NC approaches such as NC matter distributions kuniyal ; bhar ; rahaman ; NCLD2 , NC gauge theory of gravity via deformed geometry NCGTG2 ; NCGTG3 , or deformed mass in gauge theory NCGTGmassdeformed1 ; NCGTGmassdeformed2 ; NCGTGmassdeformed3 , which did not predict the existence of a stable photon sphere near the event horizon.